Your company provides a variety of delivery services. Management wants to know the volume of a particular delivery that would generate $10,300 per month in operating profits before taxes. The company charges $20 per delivery.

The controller’s office has estimated overhead costs at $9,000 per month for fixed costs and $12 per delivery for variable costs. You believe that the company should use regression analysis. Your analysis shows the results to be:

Monthly overhead = $24,942 + $10.70 per delivery

Your estimate was based on the following data.

Month Overhead Costs Number of Deliveries
1 $142,730 11,580
2 151,820 12,330
3 192,740 15,810
4 141,080 11,400
5 203,790 12,930
6 180,600 14,880
7 159,720 12,660
8 183,960 15,210
9 194,350 15,600
10 150,320 12,120
11 154,280 12,780
12 185,010 15,450
13 183,280 14,730

The company controller is somewhat surprised that the cost estimates are so different. You have been asked to recheck your work and see if you can figure out the difference between your results and the controller’s results.

Determine the amount of deliveries needed to earn operating profits of $10,300. (Round your intermediate calculations to 2 decimal places. Roundup "unit" answer to the nearest whole number.)
Answer #1

Amount of deliveries needed = (fixed cost + desired operating profit)/ contribution margin

Contribution margin = 20-10.70 = 9.30

Amount of deliveries needed = (24942+10300)/9.30 = 3789 deliveries